If x+y=K is normal to y^2= 12x ,then `k` is (A)3 (B)9(C)-9(D)-3

Grade:12th pass

1 Answers

Rabin Kar
6 years ago
B(9)comparing with the standard parabola normal at `t` is y+tx=6t+3t^3.comparing with the given eqn x+y=k, t/1=1=(6t+3t^3)/k=>k=9